Calculate the cell potential for the following reaction as written at 25.00 °C, given that [Zn2+ ] = 0.850 M and [Ni2+ ] = 0.0130 M. Standard reduction potentials can be found here. Zn(s) + Ni2+(aq)\rightleftharpoonsZn2+(aq) + Ni(s)
Answer #1

In the reaction, Zn is oxidising and Ni is reducing.

Reduction potential of Zn is -0.762 V

Oxidation potential of Zn will be +0.762 V.

Reduction potential of Ni is -0.236 V

Eo = Eo(Zn) + Eo(Ni)

= 0.762 - 0.236 = 0.526 V

[Zn2+ ] = 0.850 M

[Ni2+ ] = 0.0130 M

n = number of electrons = 2

At temperature = 25 oC, EMF of cell = Eo - 0.06/n log [Zn2+]/[Ni2+]

= 0.526 - 0.06/2 log[0.850/0.0130]

= 0.526 - 0.03 log [65.38]

= 0.526 - 0.03*1.815

= 0.526 - 0.054

= 0.472 V

cell potential = 0.472 V